OpenStudy (anonymous):

The equation of Circle P is ( x − 3)2 + (y + 2)2 = 34 and the equation of circle Q is (x − 4)2 + (y + 5)2 = 52. What is the distance between the centers of these circles? The equation of Circle P is ( x − 3)2 + (y + 2)2 = 34 and the equation of circle Q is (x − 4)2 + (y + 5)2 = 52. What is the distance between the centers of these circles? @Mathematics

OpenStudy (anonymous):

Center of Circle P is (3,-2) and the Center of Circle Q is (4,-5) And the distance formula is \[\sqrt{(x2 - x1)^2 + (y2 - y1)^2}\] Just plug in the numbers and simplify.

OpenStudy (anonymous):

Wow thanks

OpenStudy (anonymous):

ive seen this somewhere before so i should be able to get it

OpenStudy (anonymous):

Its pretty simple. Do you know how i got the centers of the circles?

OpenStudy (anonymous):

no how?

OpenStudy (anonymous):

Ok, the standard equation of a circle is \[(x-h)^2 + (y-k)^2 = r^2\] Where the center is at \[(h,k)\]

OpenStudy (anonymous):

(0,0) ?

OpenStudy (anonymous):

So your equation looks like this \[(x-3)^2 + (y+2)^2 = 34\] So comparing this with the standard equation of a circle, what would h and k be?

OpenStudy (anonymous):

i cant figure it out

OpenStudy (anonymous):

\[(x-h)^2 + (y-k)^2 = r^2 = standard\]\[(x-3)^2 + (y+2)^2 = 34 = yours\]

OpenStudy (anonymous):

They are exactly the same format.. all you do is find h in the standard and then follow it down to your problem.

OpenStudy (anonymous):

so h would be 3 and k would be -2

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

the tricky part is the -2.. since the standard equation reads \[(y-k)^2\] but you have \[(y+2)^2\] so in order to get that you subtract a negative like this \[(y-(-2))^2\] subtracting a negative is like adding so you end up with \[(y+2)^2\]

OpenStudy (anonymous):

yeah its like multiplaing a negative by a negative in a sense so it remains positive.

OpenStudy (anonymous):

multiplying i mean

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So you think you understand how to find the center?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

thanks for the help